# Tensor Decomposition Networks for Fast Machine Learning Interatomic Potential Computations

### v1

@@ 5 Experiments
The effectiveness of our method is assessed on two benchmarks: the PubChemQCR dataset, which contains millions of molecular-dynamics snapshots, and the established Open Catalyst Project (OCP) datasets. Section 5.1 describes both benchmarks and model configurations. Section 5.2 then analyses the computational efficiency of our approach as the maximum angular degree in the tensor product is varied. Section 5.3 compares our model with several baselines on PubChemQCR and its subset PubChemQCR-S, while Section 5.4 and Section 5.5 report results of our model compared to several tensor-product baselines on OC20 and OC22 from OCP, respectively.
5.1 Experimental Setting
Table 1: Comparison of model performance on energy and force predictions for the PubChemQCR and the PubChemQCR-S dataset. Our model is trained to compare against several baseline methods on the PubChemQCR-S dataset, including SchNet (schutt2018schnet, ), PaiNN (schutt2021equivariant, ), MACE (batatia2022mace, ), PACE (xu2024equivariant, ), FAENet (duval2023faenet, ), NequIP (batzner20223, ), SevenNet (park2024scalable, ), Allegro (musaelian2023learning, ), and Equiformer (liao2022equiformer, ), and. On the full PubChemQCR dataset, we compare performance with SchNet (schutt2018schnet, ) and PaiNN (schutt2021equivariant, ).
@@ 5.1 Experimental Setting
Dataset Model Energy MAE Force MAE Energy MAE Force MAE
(meV) $\downarrow$ (meV/Å) $\downarrow$ (meV) $\downarrow$ (meV/Å) $\downarrow$
@@
The second dataset used in this study is the Open Catalyst Project (OCP) (zitnick2020introduction, ), which publishes extensive open datasets of DFT relaxations for adsorbate–catalyst surfaces and hosts public leaderboard challenges. The flagship releases, OC20 (chanussot2021open, ) and OC22 (tran2023open, ), encompass thousands of chemical compositions, crystal facets, and adsorbates, serving as comprehensive benchmarks for surrogate models aiming to replace computationally intensive quantum-mechanical calculations. Each release defines multiple tasks, including Initial-Structure-to-Relaxed-Energy (IS2RE) and Structure-to-Energy-and-Force (S2EF), which require accurate predictions of total energies, per-atom forces, and relaxed geometries. We conduct experiments on OC20 IS2RE and S2EF, as well as OC22 IS2RE, using the data splits and configurations specified in the official OCP repository.
@@
Figure 3: Scaling behaviours of the CP-decomposition-based tensor product under different maximum angular degree scheduler: (a) approximation error, (b) $\rm{SO}(3)$-equivariance error, and (c) runtime.
@@
We approximate the fully connected CG tensor product using the rank-$R$ CP decomposition across various configurations of the maximum angular degree $L$. Four rank schedules are evaluated: $R=(L+1)^{4}$, $R=16\log^{2}(L+1)$, $R=7L$, and $R=7L^{2}$. For each configuration we measure (i) Approximation Error: The relative CP tensor product error, calculated as $\lVert{\bm{M}}({\bm{x}}\otimes{\bm{y}})-\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert_{F}/\lVert{\bm{M}}({\bm{x}}\otimes{\bm{y}})\rVert_{F}$; (ii) Equivariance Error: The expected $\mathrm{SO}(3)$-equivariance error, defined as $\mathbb{E}_{{\bm{R}},{\bm{x}},{\bm{y}}}\bigl{[}\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\bigr{]}$, where the expectation is averaged over $1000$ random rotations ${\bm{R}}$ and random vectors ${\bm{x}},{\bm{y}}$; and (iii) Execution Time: The runtime for the CP-decomposition-based tensor product under each rank schedule. Our results demonstrate that the logarithmic schedule $R=7L^{2}$ consistently achieves the lowest approximation and equivariance errors across all $L$, while requiring significantly less computation time compared to the CG tensor product baseline. Thus, this schedule provides the best balance between accuracy and computational cost and is adopted for all subsequent experiments.
@@
Building on the capabilities of graph transformers, we develop the Tensor-Decomposition Network (TDN) by modifying the publicly available Equiformer architecture (liao2022equiformer, ). Specifically, we replace every channel-wise linear projection, normalization layer, and activation in Equiformer with batched matrix operations that act simultaneously on the multiplicity dimension of each irrep block, eliminating the need for costly slicing and reshaping between tensor and vector representations. Additionally, the core CG tensor products within the self-attention mechanism are substituted with our rank-$R$ CP-decomposition-based tensor product from Section 4.1, integrated with the path-weight-sharing scheme from Section 4.2. This design preserves the expressive power of Equiformer while significantly improving memory and computational efficiency. The model and training configurations for TDN are described in Appendix C.
@@
For the PubChemQCR benchmark, we reimplement each baseline model based on its official repository. Hyperparameters are adopted from the best configurations reported in the original papers or, if unspecified, are tuned using the PubChemQCR-S dataset. All model and training configurations of baseline models are described in Appendix C.
5.2 Efficiency of CP-Decomposition-Based Tensor Product
Model Throughput (samples/sec) Parameter
Equiformer (L=1) 470.9 8.86M
TDN (L=1) 583.1 4.42M
Equiformer (L=2) 122.5 18.19M
TDN (L=2) 199.7 4.42M
Equiformer (L=3) 48.4 33.43M
TDN (L=3) 142.9 4.42M
To evaluate the speed-up achieved by our CP-decomposition-based tensor product in Section 4.1, we benchmark its wall-clock runtime against the fully connected CG tensor product implementation in e3nn. As shown in Fig. 3(c), the proposed approximation accelerates by factors of 4.0×,4.8×,15.0×,26.7×,47.6×,4.0$\times$,4.8$\times$,15.0$\times$,26.7$\times$,47.6$\times$,4.0 × , 4.8 × , 15.0 × , 26.7 × , 47.6 × , and 83.6×83.6$\times$83.6 × for maximum degrees $L=1,2,3,4,5,$ and $6$, respectively.
Since TDN is derived from Equiformer by replacing each CG block with its CP decomposition and incorporating path-weight sharing, we further evaluate end-to-end throughput and parameter count for both networks under identical irrep configurations across varying $L$. All experiments are conducted on a single NVIDIA A100-80GB GPU with batch size 128 and identical input irreps. The results in Table 2 show that TDN processes 1.23×1.23$\times$1.23 × to 2.95×2.95$\times$2.95 × more structures per second and uses $50\%$$87\%$ fewer parameters than Equiformer. As the maximum degree $L$ increases, TDN achieves higher throughput and requires fewer parameters.
@@ 5.3 Results on PubChemQCR
We evaluate our method on both PubChemQCR-S and the full PubChemQCR dataset. For the small split we compare against nine state-of-the-art models: SchNet (schutt2018schnet, ), PaiNN (schutt2021equivariant, ), MACE (batatia2022mace, ), PACE (xu2024equivariant, ), FAENet (duval2023faenet, ), NequIP (batzner20223, ), SevenNet (park2024scalable, ), Allegro (musaelian2023learning, ), and Equiformer (liao2022equiformer, ). On the full dataset we benchmark against SchNet and PaiNN, the only baselines that scale to its size within our hardware budget. Performance is reported as mean absolute error (MAE) for energies and root-mean-square error (RMSE) for forces. The results are shown in Table 1, the proposed TDN model achieves the lowest energy and force prediction errors across PubChemQCR-S and the full PubChemQCR dataset, significantly outperforming all baseline methods. Moreover, the performance further improves as the dataset size increases.
5.4 Results on OC20
Table summarizes our performance on the two principal OC20 validation tasks: IS2RE-Direct, which predicts the relaxed adsorption energy directly from the initial geometry (no noisy-node auxiliary loss), and S2EF on the 2M subset, which jointly regresses per-atom forces and total energies.
Table 3: Comparison of model performance on energy and force predictions for OC20 S2EF 2M validation set. Our model is trained to compare against several baseline methods, including SchNet (schutt2018schnet, ), DimeNet++ (gasteiger2020fast, ), SpinConv (shuaibi2021rotation, ), GemNet-OC (gasteiger2022gemnet, ), eSCN (passaro2023reducing, ), and EquiformerV2 (liao2024equiformerv, ).
Model Energy MAE Force MAE
(meV) (meV/Å)
SchNet 1400 78.3
DimeNet++ 805 65.7
SpinConv 406 36.2
GemNet-OC 286 25.7
eSCN 283 20.5
EquiformerV2 285 20.5
TDN 257 30.8
We benchmark against the strongest tensor-product architectures reported to date, including Equiformer (liao2022equiformer, ), Equiformer-v2 (liao2024equiformerv, ), and eSCN (passaro2023reducing, ). Metrics follow the official OC20 protocol: energy mean absolute error (MAE, eV) and energy within threshold (EwT, %) in IS2RE-Direct for four validation sub-splits: distribution adsorbates and catalysts (ID), out-of-distribution adsorbates (OOD-Ads), out-of-distribution catalysts (OOD-Cat), and out-of-distribution adsorbates and catalysts (OOD-Both), and energies MAE (meV) and forces MAE (meV/Å) in S2EF 2M validation set. The S2EF results are shown in Table 3. Our TDN model achieves superior performance in energy prediction while maintaining comparable accuracy in force prediction. The IS2RE-Direct results are presented in Table 4, demonstrating that our model achieves performance comparable to Equiformer while being more efficient, as detailed in Section 5.2. This highlights the effectiveness of our architecture in maintaining expressivity while significantly reducing computational costs.
Both the IS2RE-Direct and S2EF 2M TDN models are trained with the optimization scheme of Equiformer (liao2022equiformer, ): AdamW with learning-rate $2\times 10^{-4}$ and weight decay $10^{-3}$, a batch size of 32, and a cosine decay schedule with 2 warm-up epochs and 0.2 warm-up factor over 30 training epochs.
Table 4: Comparison of model performance on energy predictions for OC20 IS2RE-Direct validation set without noisy-node auxiliary loss. Our model is trained to compare against several baseline methods, including SchNet (schutt2018schnet, ), DimeNet++ (gasteiger2020fast, ), GemNet-dT (gasteiger2021gemnet, ), SphereNet (liu2022spherical, ), Equiformer (liao2022equiformer, ) and EquiformerV2 (liao2024equiformerv, ).
@@ 5.5 Results on OC22
Table 5: Comparison of model performance on energy predictions for OC22 IS2RE-Direct testing set. We compare with several baseline methods, including SchNet (schutt2018schnet, ), DimeNet++ (gasteiger2020fast, ), PaiNN (schutt2021equivariant, ), GemNet-dT (gasteiger2021gemnet, ), and coGN (ruff2302connectivity, ).
@@
Following the OC20 protocol, we assess our model on the OC22 IS2RE-Direct test set, in which the energy is predicted directly from the initial structure without the noisy-node auxiliary loss. Metrics include the average, in-distribution (ID) and out-of-distribution (OOD) mean absolute error (MAE, eV) obtained from the four standard OC22 splits with the same split protocol as mentioned in OC20. For training, the TDN model is trained with the same optimization scheme as for OC20 IS2RE-Direct but for 1000 epochs. As shown in Table 5, TDN achieves the lowest in-distribution and average MAE and second best performance on out-of-distribution MAE, validating the high effectiveness of TDN.

### v2

@@ 4 Experiments
The effectiveness of our method is assessed on two benchmarks: the PubChemQCR dataset, which contains millions of molecular‑dynamics snapshots, and the established Open Catalyst Project (OCP) datasets. Section˜4.1 describes both benchmarks and model configurations. Section˜4.2 compares our model with several baselines on PubChemQCR and its subset PubChemQCR-S, while Section˜4.3 and Section˜4.4 report results of our model compared to several tensor‑product baselines on OC20 and OC22 from OCP, respectively. Section˜4.5 then analyses the computational efficiency of our approach as the maximum angular degree in the tensor product is varied. Section˜4.6 further reports ablations of the proposed components across our model and a second widely used architecture.
4.1 Experimental Setup
Table 1: Comparison of model performance on energy and force predictions for the PubChemQCR and the PubChemQCR-S dataset. Our model is trained to compare against several baseline methods on the PubChemQCR-S dataset, including SchNet (schutt2018schnet, ), PaiNN (schutt2021equivariant, ), MACE (batatia2022mace, ), PACE (xu2024equivariant, ), FAENet (duval2023faenet, ), NequIP (batzner20223, ), SevenNet (park2024scalable, ), Allegro (musaelian2023learning, ), and Equiformer (liao2022equiformer, ). On the full PubChemQCR dataset, we compare our model with SchNet (schutt2018schnet, ) and PaiNN (schutt2021equivariant, ). The best results are shown in bold.
@@ 4.1 Experimental Setup
Dataset Model Energy MAE Force RMSE Energy MAE Force RMSE
(meV/atom) $\downarrow$ (meV/Å) $\downarrow$ (meV/atom) $\downarrow$ (meV/Å) $\downarrow$
@@
The second dataset used in this study is the Open Catalyst Project (OCP) (zitnick2020introduction, ), which publishes extensive open datasets of DFT relaxations for adsorbate–catalyst surfaces and hosts public leaderboard challenges. The flagship releases, OC20 (chanussot2021open, ) and OC22 (tran2023open, ), encompass thousands of chemical compositions, crystal facets, and adsorbates, serving as comprehensive benchmarks for surrogate models aiming to replace computationally intensive calculations. Each release defines multiple tasks, including Initial-Structure-to-Relaxed-Energy (IS2RE), which require accurate predictions of total energies, per-atom forces, and relaxed geometries. We conduct experiments on OC20 IS2RE and OC22 IS2RE, using the data splits and configurations specified in the official OCP repository.
@@
Figure 3: Scaling behaviours of the CP-decomposition-based tensor product under different maximum angular degree scheduler: (a) approximation error, (b) $\rm{SO}(3)$-equivariance error, and (c) runtime. The error and runtime of the CP decomposition-based tensor product depend on the chosen rank, multiplicities, and the maximum angular degree, and are not tied to a specific dataset.
@@
We approximate the fully connected CG tensor product using the rank-$R$ CP decomposition across various configurations of the maximum angular degree $L$. Four rank schedules are evaluated: $R=(L+1)^{4}$, $R=16\log^{2}(L+1)$, $R=7L$, and $R=7L^{2}$. For each configuration we measure (i) Approximation Error: The relative CP tensor product error, calculated as $\lVert{\bm{M}}({\bm{x}}\otimes{\bm{y}})-\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert_{F}/\lVert{\bm{M}}({\bm{x}}\otimes{\bm{y}})\rVert_{F}$; (ii) Equivariance Error: The expected $\mathrm{SO}(3)$-equivariance error, defined as $\mathbb{E}_{{\bm{R}},{\bm{x}},{\bm{y}}}\bigl[\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\bigr]$, where the expectation is averaged over $1000$ random rotations $\bm{R}$} and random vectors ${\bm{x}},{\bm{y}}$; and (iii) Execution Time: The runtime for the CP-decomposition-based tensor product under each rank schedule. Our results demonstrate that the logarithmic schedule $R=7L^{2}$ consistently achieves the lowest approximation and equivariance errors across all $L$, while requiring significantly less computation time compared to the CG tensor product baseline. Thus, this schedule provides the best balance between accuracy and computational cost and is adopted for all subsequent experiments. Further rank ablations are described in Appendix˜D.
@@
Building on the capabilities of graph transformers, we develop the Tensor-Decomposition Network (TDN) by modifying the publicly available Equiformer architecture (liao2022equiformer, ). Specifically, we replace every channel-wise linear projection, normalization layer, and activation in Equiformer with batched matrix operations that act simultaneously on the multiplicity dimension of each irrep block, eliminating the need for costly slicing and reshaping between tensor and vector representations. Additionally, the depth-wise tensor product mechanism is removed and the core CG tensor products within the self-attention mechanism are substituted with our rank-$R$ CP-decomposition-based tensor product from Section˜3.1, integrated with the path-weight-sharing scheme from Section˜3.2. This design preserves the expressive power of CG tensor product while significantly improving memory and computational efficiency. We further integrate the equivariant linear layer with the path-weight-sharing scheme. For the OCP dataset, the same model architecture is applied and a subset of experiments additionally incorporate initial node embeddings following (qu2024the, ). The detailed model configurations for TDN are described in Appendix˜C.
@@
For the PubChemQCR benchmark, we reimplement each baseline model based on its official repository. Hyperparameters are adopted from the best configurations reported in the original papers or, if unspecified, are tuned using the PubChemQCR-S dataset. All model configurations of baseline models are described in Appendix˜C.
@@ 4.2 Results on PubChemQCR
We evaluate our method on both PubChemQCR‑S and the full PubChemQCR dataset. For the small split we compare against nine state‑of‑the‑art models: SchNet (schutt2018schnet, ), PaiNN (schutt2021equivariant, ), MACE (batatia2022mace, ), PACE (xu2024equivariant, ), FAENet (duval2023faenet, ), NequIP (batzner20223, ), SevenNet (park2024scalable, ), Allegro (musaelian2023learning, ), and Equiformer (liao2022equiformer, ). On the full PubChemQCR dataset, we benchmark against SchNet and PaiNN, the only baselines that scale to its size within our hardware budget. Performance is reported as mean absolute error (MAE) for energies and root‑mean‑square error (RMSE) for forces over the validation and testing splits. The results are shown in Table˜1, the proposed TDN model achieves the lowest energy and force prediction errors across PubChemQCR‑S and the full PubChemQCR dataset, outperforming all baseline methods. In addition, the performance of TDN improves further as the size of the dataset increases.
Training Setup. Across both the PubChemQCR and PubChemQCR-S benchmarks, we adopt a uniform training protocol: a cutoff radius of 4.5 Å; the Adam optimizer with an initial learning rate of $5\times 10^{-4}$; and a ReduceLROnPlateau scheduler with a patience of 2 epochs. All models are trained for up to 100 epochs on PubChemQCR-S and up to 15 epochs on the full PubChemQCR dataset. Unless otherwise noted, experiments are executed on NVIDIA A100-80GB GPUs.
Table 2: Comparison of model performance on energy predictions for OC20 IS2RE-Direct validation set without noisy-node auxiliary loss. Our model is trained to compare against several baseline methods, including SchNet (schutt2018schnet, ), DimeNet++ (gasteiger2020fast, ), GemNet-dT (gasteiger2021gemnet, ), SphereNet (liu2022spherical, ), Equiformer (liao2022equiformer, ) and EquiformerV2 (liao2024equiformerv, ). The best results are shown in bold and the second best results are shown with underlines.
@@
4.3 Results on OC20
Table˜2 summarizes our performance on the principal OC20 task IS2RE‑Direct, which predicts the relaxed adsorption energy directly from the initial geometry (no noisy‑node auxiliary loss). We benchmark against the widely-used baselines reported to date, including SchNet (schutt2018schnet, ), DimeNet++ (gasteiger2020fast, ), GemNet-dT (gasteiger2021gemnet, ), SphereNet (liu2022spherical, ), Equiformer (liao2022equiformer, ) and EquiformerV2 (liao2024equiformerv, ). Metrics follow the official OC20 protocol: energy mean absolute error (MAE, eV) and energy within threshold (EwT, %) in IS2RE-Direct for four validation sub-splits: distribution adsorbates and catalysts (ID), out-of-distribution adsorbates (OOD-Ads), out-of-distribution catalysts (OOD-Cat), and out-of-distribution adsorbates and catalysts (OOD-Both). The IS2RE‑Direct results are presented in Table˜2, demonstrating that our model achieves performance comparable to Equiformer while being more efficient, as detailed in Section˜4.5. This highlights the effectiveness of our architecture in maintaining accuracy while significantly reducing computational costs.
Training Setup. For the IS2RE-Direct task, we follow Equiformer’s optimization setup (liao2022equiformer, ) by an AdamW optimizer with a learning rate of $2\times 10^{-4}$ and a weight decay of $10^{-3}$, a batch size of 32, and a cosine-decay learning-rate schedule. A warm-up is employed for 2 epochs on IS2RE-Direct, with a warm-up factor of 0.2; the cosine decay then runs over 30 training epochs. IS2RE-Direct experiments are run on a single NVIDIA RTX A6000-48GB GPU.
@@ 4.4 Results on OC22
Table 3: Comparison of model performance on energy predictions for OC22 IS2RE-Direct testing set. We compare with several baseline methods, including SchNet (schutt2018schnet, ), DimeNet++ (gasteiger2020fast, ), PaiNN (schutt2021equivariant, ), GemNet-dT (gasteiger2021gemnet, ), and coGN (ruff2302connectivity, ). The best results are shown in bold and the second best results are shown with underlines.
@@
Following the OC20 evaluation protocol, we evaluate on the OC22 IS2RE-Direct test set, which predicts relaxed energy directly from the initial structure and omits the noisy-node auxiliary loss. We benchmark against strong baselines up to date as listed in Table˜3 and report mean absolute error (MAE, eV) for the in-distribution (ID) and out-of-distribution (OOD) splits by averaging across the four standard OC22 sub-splits using the same split scheme as OC20 IS2RE-Direct task. As summarized in Table˜3, TDN achieves the lowest ID MAE and the best overall average MAE, and it attains the second-best OOD MAE, validating the high effectiveness of TDN.
Training Setup. On OC22 IS2RE, we use the same optimization setup as OC20 by an AdamW optimizer with a learning rate of $2\times 10^{-4}$, a weight decay of $10^{-3}$, a batch size of 32, and a cosine learning-rate decay with a 2-epoch warm-up for 1000 epochs on a single NVIDIA A100-80GB GPU.
4.5 Efficiency of CP-Decomposition-Based Tensor Product
Model Throughput (samples/sec) Parameter
Equiformer (L=1) 311.7 12.1M
TDN (L=1) 770.8 ($\times$ 2.47) 4.5M ($\times$ 0.37)
Equiformer (L=2) 71.9 27.9M
TDN (L=2) 312.4 ($\times$ 4.34) 4.5M ($\times$ 0.16)
Equiformer (L=3) 26.1 54.7M
TDN (L=3) 220.4 ($\times$ 8.44) 4.5M ($\times$ 0.08)
To evaluate the speed-up achieved by our CP-decomposition-based tensor product in Section˜3.1, we benchmark its inference runtime against the fully connected CG tensor product implementation in e3nn. As shown in Fig.˜3(c), the proposed approximation accelerates by factors of $4.0\times$, $4.8\times$, $15.0\times$, $26.7\times$, $47.6\times$, and $83.6\times$ for maximum degrees $L=1,2,3,4,5,$ and $6$, respectively.
Since TDN is derived from Equiformer by replacing each CG block with CP decomposition and incorporating path-weight sharing, and removing depth-wise tensor product mechanism, we further benchmark end-to-end throughput and parameter count for both networks on a single NVIDIA A100-80GB GPU and Xeon Gold 6258R processor with a batch size of 128; detailed model configurations are provided in Appendix˜D. The results in Table˜4 show that TDN processes $2.47\times$ to $8.44\times$ more structures per second and uses $63\%$$92\%$ fewer parameters than Equiformer. As the maximum degree $L$ increases, TDN achieves higher throughput and requires fewer parameters. A comprehensive time ablation of CP decomposition and path-weight sharing mechanism is provided in Table˜8.
4.6 Ablation Study of CP-Decomposition-Based Tensor Product
Table 5: Ablation study of the path-weight sharing tensor product (PS) and CP-decomposition-based tensor product (CP). Results are shown for TDN and NequIP (batzner20223, ) on PubChemQCR-S dataset.
Model Validation
Energy MAE Force RMSE Training Time
(eV) $\downarrow$ (eV/Å) $\downarrow$ (min/epoch) $\downarrow$
TDN w/o CP + PS 0.01274 0.09274 19.0
TDN 0.01295 0.09653 4.2 ($\times$ 0.22)
NequIP 0.01122 0.09009 7.5
NequIP + CP + PS 0.01115 0.09249 2.0 ($\times$ 0.26)
To further evaluate the performance and efficiency contributions of our proposed component, we perform ablations on the PubChemQCR-S dataset by removing the path-weight sharing tensor product and the CP decomposition of TDN while holding all other components fixed. As shown in Table˜5, starting from TDN, disabling both mechanisms yields slightly lower validation energy MAE and force RMSE while reducing training time roughly 78%, indicating only a minor effect on predictive accuracy. TDN is trained with a maximum degree $L=2$, four graph transformer layers, a hidden dimension of 64, and for 60 epochs. We also evaluate NequIP (batzner20223, ) by augmenting the base model with path-weight sharing tensor product and CP decomposition. This preserves downstream accuracy while cutting computational cost nearly 74%. NequIP is trained for 100 epochs with a maximum degree $L=2$, four interaction blocks, and multiplicity 64. All experiments are conducted on a single NVIDIA RTX A6000-48GB GPU.
5 Limitations
Our approach accelerates $\rm{SO}(3)$‑equivariant tensor products by replacing the full CG tensor product with a low‑rank CP decomposition. We substantiate both its theoretical speed‑up and its empirical efficacy across multiple benchmarks. One limitation of our approach is rank selection: the minimal CP rank needed to attain a given approximation error is unknown in general, and computing it exactly is NP-hard (kolda2009tensor, ). We therefore employ an empirical rank scheduler; however, this scheduler is tailored to the CG tensor and must be re-derived when extending to other equivariant tensor products. A second limitation is that path-weight sharing mechanism, while reducing parameters and memory, can introduce a small accuracy drop. In future work, we will (i) study broader families of group-equivariant tensor products, e.g., $\rm{SO}(2)$ convolution (passaro2023reducing, ) and Gaunt tensor products (luoenabling, ), to characterize their optimal rank profiles and develop general rank-selection criteria and adaptive rank selector, and (ii) design adaptive path-selection and grouping strategies that preserve the efficiency benefits of weight sharing while recovering performance.